OpenStudy (anonymous):

Log(10) (root 10)= ?

OpenStudy (anonymous):

\[\log_{10}\sqrt{10}\] \]

OpenStudy (anonymous):

1/2

OpenStudy (anonymous):

But how do I get power with a root?

OpenStudy (anonymous):

10^?=\[\sqrt{10}\]?

OpenStudy (anonymous):

root(10)=10^(1/2)